L {t^{2}*H'(t-3)}


1
Expert's answer
2021-08-09T08:04:39-0400

"L = ( {t^{2}(t-3)})\\
t-3 =L;\quad t²=1\\
t-3 =L;\quad t=\sqrt{1² },=1\\
t-3 =L;\quad t=1\\
1-3 =L;\quad t=1\\
L =-2; \quad \ \ \ t=1"